# Magnetic field measurements in a sample of Class I and Flat-Spectrum protostars observed with SPIRou

### 18 Mar 2026 v1

@@ Abstract
Context. Magnetic fields play a crucial role throughout stellar evolution, regulating angular momentum, channelling accretion, and launching jets and outflows. While the magnetic properties of Classical T Tauri Stars (CTTS) have been extensively characterised, those of their progenitors, Class I and Flat-Spectrum (FS) protostars, remain poorly constrained due to their embedded nature, which provides an observational challenge.
Aims. We aim to detect and characterise the large-scale magnetic fields in a sample of Class I and FS protostars. These young stars are expected to host strong magnetic fields generated by dynamo processes in their largely convective interiors.
Methods. We have used SPIRou, a high-resolution spectropolarimeter working in the near-infrared domain, to analyse the polarised light of Class I and FS protostars. We used the Least Squares Deconvolution (LSD) technique to perform the magnetic analysis and measure the longitudinal magnetic fields from circularly polarised Stokes $V$ profiles.
Results. We report new large-scale magnetic field detections in 5 FS protostars. Including the previous detection of the large-scale magnetic-field in the V347 Aur FS-protostar, 40% of our final sample of 15 protostars is confirmed to be magnetic. These magnetic stars show clear Zeeman signatures, with longitudinal field strengths ranging from $\sim$80 to $\sim$200 G in absolute value. The remaining stars exhibit no detectable Stokes $V$ signature, but the estimated upper limits on a hidden dipolar field range from 500 G to more than 5 kG. For stars in which no magnetic fields are detected, it is still conceivable that a magnetic field exists, but is either intrinsically weak, highly complex and dominated by small-scale structures, or canceled out in integrated spectropolarimetric signals due to opposing polarities.
Conclusions. We show that Class I and FS protostars can host large-scale magnetic fields with strengths that may be weaker than in more evolved CTTS. This supports the idea that magnetic processes are already active during the main accretion phase and may influence star-disk interactions from the earliest stages.
@@ 1 Introduction
Star formation occurs in dense cores within molecular clouds. These clouds are magnetised, and the cores, called pre-stellar cores (Ward-Thompson et al., 1994), collapse under gravity, eventually giving birth to protostars. During this early evolution, magnetic fields are thought to play a significant role in several processes, such as accretion, ejection, and interaction with the circumstellar environment (Lada, 1987). Through their role in these processes that significantly impact stellar rotation and evolution, magnetic fields shape the close environment of the protostar (Bouvier et al., 2007; Vaytet et al., 2018).
@@
Regulation of angular momentum is made through magnetospheric accretion and ejection processes during the pre-main-sequence (PMS) phase (Bouvier et al., 2014; Lee et al., 2017). In Classical T Tauri Stars (CTTS), the presence of large-scale magnetic fields is well established (Donati et al., 2011). These fields, typically of kilo gauss strength, are responsible for truncating the inner disk and channelling the accreted material onto the stellar surface (Bouvier et al., 2007, 2014; Hartmann et al., 2016). The magnetic properties of CTTS (between 0.5-1.3 $M_{\odot}$) are characterised by a dominant poloidal field, with an average strength between 1 and 2 kG (Zaire et al., 2024).
@@
While magnetism and its impact on the star-disk interaction have been densely explored during the PMS phase (Romanova et al., 2015), our current knowledge on the magnetic properties, and on the star-disk interaction in earlier phases remains limited. During those phases, the protostar experiences intense accretion, often through bursts, and grows in mass until it approaches its final mass (Baraffe et al., 2009; Fischer et al., 2023). Such major internal structural changes may have a strong impact on the stellar magnetic field, which in turn may influence the connection between the star and its accretion disk (Kunitomo et al., 2017). The progenitors of the CTTS (Class II) are the Class I protostars. These objects are still embedded within their envelopes of gas and dust (Lada, 1987). The protostars transiting from Class I to Class II are observed as Flat-Spectrum (FS) sources (Greene et al., 1994). Due to their dense and obscuring environments, direct observation of Class I protostars and detection of photospheric features is challenging (Connelley and Greene, 2010; Doppmann et al., 2005), making it difficult to study their surfaces and, therefore, their magnetic fields.
@@
For several reasons, we expect Class I protostars to be strongly magnetised, as in CTTS. They have evolved enough to set off convection in their full interior (Palla, 1998; Baraffe et al., 2017) and thus are capable of hosting strong magnetic fields generated by dynamo processes. In other types of fully convective stars, such as the M dwarfs on the main-sequence, strong dynamo magnetic fields are routinely observed (Donati et al., 2008; Morin et al., 2010; Donati et al., 2023). Signs of magnetospheric accretion, as in the CTTS, have been detected in Class I sources (Doppmann et al., 2005; Fiorellino et al., 2023), suggesting the presence of strong large-scale magnetic fields at the surface of the embedded star. Powerful flares have been detected in X-ray observations of Class 0 and Class I protostars (Grosso et al., 1997, 2020; Montmerle et al., 2000), similar to those present in the Sun or more evolved stars but much more energetic (Feigelson et al., 2007). These emissions are clues of intense magnetic activity. These studies suggest, therefore, that strong magnetic fields may already be in place during the Class I phase and regulate mass inflow and angular momentum loss early on.
@@
A recent study by Flores et al. (2024) has investigated a sample of 42 Class I and FS protostars, and measured their total surface magnetic field from the Zeeman broadening of spectral lines. They found a median magnetic field strength of 1.9 kG, which is consistent with those found in Class II sources and with earlier studies of Class I protostars (Johns-Krull et al., 2009), indicating that Class I and FS sources are indeed strongly magnetic. However, this study cannot inform us on the large-scale components of their magnetic fields, which interact with the disc. To understand whether magnetospheric accretion and ejection processes are in place at that age, and are potentially at the origin of the jets observed in almost all Class I, we need to characterise their large-scale magnetic field.
@@
Furthermore, dynamo processes are not well understood, especially in fully convective stars. The imprint of the dynamo processes in stellar evolution is still an open question (Moss, 2003; Donati et al., 2008). In particular, simulations have shown that the initial conditions at the launch of dynamo may impact the future properties of dynamo magnetic fields, and may explain the bistability of magnetic fields observed in fully convective M dwarfs (Gastine et al., 2012, 2013; Raynaud et al., 2015). While the magnetic energy produced by dynamo processes is an important characteristic to constrain the models, another factor is the ratio of poloidal to toroidal fields, which is only accessible by measuring the magnetic geometry.
@@
The topology and strength of the large-scale magnetic field can be obtained with the Zeeman Doppler Imaging (ZDI) technique (Semel, 1989). This technique uses time-series of circularly polarised and intensity spectra well sampled over the rotation period of the star using a high spectral resolution spectropolarimeter. By assuming the magnetic field components can be modelled as spherical harmonics, and by using an inversion technique, ZDI provides us with the maps of the radial, azimuthal, and meridional components of the field (Donati et al., 2006; Kochukhov, 2016). This technique has been largely employed to map the magnetic field of various types of stars, including CTTS (e.g. Donati and Gregory, 2009; Zaire et al., 2024).
Up to now, ZDI has never been largely performed in Class I and FS protostars.
@@
Recent advances in instrumentation have opened new possibilities for investigating these elusive protostars. Among these instruments is SPIRou (SPectropolarimètre InfraRouge), a spectropolarimeter installed on the 3.6-meter Canada-France-Hawaii Telescope (CFHT). SPIRou (Donati et al., 2020) operates in the near-infrared (nIR) domain, which allows us to pierce through their dusty envelope to probe the photosphere of Class I protostars. SPIRou’s high-resolution, high-sensitivity measurements in the $YJHK$ bands are particularly well suited to study Class I protostars, which are largely inaccessible to optical spectropolarimeters.
@@
Before performing a monitoring for each source of our sample of Class I and FS protostars, to map their magnetic fields, we performed a snapshot survey with SPIRou. Our aim is to first detect their Zeeman signature in circularly polarised (Stokes $V$) spectra. In future work, we will monitor with SPIRou only those with a Zeeman detection to perform ZDI.
In this study, we present the analysis of our snapshot sample. It is structured as follows. We describe the observations, data reduction, and spectral and polarimetric analyses in Section 2. Our results are presented in Section 3 and discussed in Section 4. We summarise our conclusions in Section 5.
@@ 2.1 Sample selection
Our goal is to characterise the magnetic fields of a sample of Class I and FS objects. To this aim, we need to detect the circular polarisation signal induced by the Zeeman effect in the photospheric lines. To build our sample, we first listed the objects in which photospheric lines had been detected in previous surveys using high-resolution optical or nIR spectrographs (White and Hillenbrand, 2004; Doppmann et al., 2005; Evans et al., 2009; Connelley and Greene, 2010; Rebull et al., 2010; Viana Almeida et al., 2012; Dunham et al., 2015). Then we first kept the sources north to a declination of -40, to ensure that they were observable from the CFHT. In a second step, we selected sources brighter than 10 magnitudes in the H band so that the total exposure time needed to detect a typical dipolar magnetic field of $\sim$0.5 kG is within 1.5 h, assuming a typical veiling in the H band of about 2 (Doppmann et al., 2005). In the star-forming regions that we explored, Class I and FS sources display H band magnitudes typically ranging from 8 to 12, depending on how deeply embedded they are or how inclined their bipolar cavity is with respect to our line of sight. As a consequence, the magnitude criterion may introduce a bias towards either the less embedded, i.e. more evolved Class I objects, or towards low inclination. Two of our targets have magnitudes H greater than 10. We had an opportunity to obtain additional observations, so we included them in our sample.
@@ 2.2 SPIRou observations and data reduction
The primary goal of SPIRou is to detect and characterise exoplanets around M-dwarfs, but also to study magnetic fields in low-mass mature and PMS (Donati et al., 2020). SPIRou covers a spectral range from 950 nm to 2 500 nm, with a resolution of 70 000 and a velocity bin per pixel of 2.3 km s-1(Donati et al., 2020). It consists of an achromatic polarimeter, which splits light into two beams associated with orthogonal states of the selected polarization, transmitted via optical fibres to an ultra-stable thermally controlled cryogenic spectrograph. The spectrograph disperses light using an echelle grating and a double-pass cross-dispersing prism train, recording the spectra on an H4RG detector.
The polarimeter consists of Fresnel rhombs and a Wollaston prism, enabling simultaneous measurements of both orthogonal states of the selected polarization (circular polarization in the present case). The operational process includes stabilising the stellar image on the entrance aperture, injecting it into optical fibres, and using calibration sources (a Fabry-Pérot interferometer, in our case) for precise wavelength and radial velocity calibration.
@@
We used V0.7.292 version of the dedicated APERO [^2] https://apero.exoplanets.ca/ pipeline (Cook et al., 2022) to reduce the observations. Compared to version V0.7.255 described in Cook et al. (2022), in the APERO versions there have been changes to the background subtraction. The thermal background spectrum is obtained by averaging the spectra acquired from telescope dark frames (i.e. with no sky scene injected). The thermal spectrum is then scaled to match the thermal emission of saturated absorption lines in the K band. An issue was noted, as persistence structures in the J band led to spurious subtraction, being mistaken for low-level thermal emission. Since one does not expect a measurable level of thermal emission below 1.9 µm (considering the rapid decline of a near-room-temperature blackbody blueward of this wavelength), the thermal emission spectrum was set to zero below 1.9 µm in the newest versions of APERO. It should be noted that, in practise, the thermal background is only subtracted within the K band; thus, this zeroing of the thermal emission at the blue end does not affect the overall thermal background correction. Finally, all individual spectra were corrected for telluric absorption (including OH lines), as described in Cook et al. (2022).
@@ 2.3 Continuum normalisation of individual spectra
To analyse the magnetic signals within the photospheric lines, a normalisation of the spectra to their continuum is needed. Automatic continuum normalisation can be performed during the extraction of the polarised spectra. However, the spectra of Class I sources have several variable emission lines that need to be taken into account on a case-by-case basis. While normalisation tools already exist, they lack specific options we aimed to incorporate (necessary for emission line spectra such as those of Class I sources). We developed a normalisation tool for SPIRou data files called Fantasio [^3] https://github.com/Lisa2626/Fantasio. The code aims to reproduce the principles of the IRAF normalisation tool called ”continuum”. The selection of the continuum points is performed by removing non-continuum points using a sigma-clipping procedure over several iterations. The normalisation is performed order by order when a 2D SPIRou spectrum is given as input. The developed interface allows users to modify normalisation parameters interactively for each order, including selecting the polynomial order of the spline and the number of spline interpolations. The sigma-clipping parameters, such as the upper and lower thresholds, and the number of iterations, can also be adjusted to optimise the fit. We allow different threshold values for points above and below the continuum to provide flexibility when both photospheric and emission lines are present within one order. Another interactive feature enables the removal of photospheric and emission lines to exclude them from the fitting process by selecting the regions to reject. This is particularly useful when dealing with spectra containing numerous strong emission lines, as in Class I sources. Normalisation can be performed manually by selecting parameters for each order, or it can be run automatically for multiple observations using pre-defined parameters.
Figure 1: Upper panel: order 32 of the spectrum of VV CrA SW without normalisation (black full line). In blue are the data points selected to fit the continuum using a sigma-clipping of 3 and 2 iterations. The spline fit of order 3 is plotted in red with a number of knots of 2. The two regions with emission lines were manually removed using the interactive tool. Lower panel: the same spectrum after normalisation. The dashed red line shows the continuum normalised to 1.
We performed a manual normalisation on both A- and B-spectra, corresponding to the two orthogonal polarisations, for one observation of each star. We then applied the automatic normalisation to all other observations of the same star, using the same fitting parameters per order as those chosen for the fit of the first observation. In general, the order of the spline fit ranges from 1 to 3, as the number of knots of the spline. The sigma clipping can vary from 1 to 3 sigma, and 3 iterations are used. An example of the normalisation process is shown in Figure 1.

### 22 Mar 2026 v2

@@ Abstract
Context. Magnetic fields play a crucial role throughout stellar evolution in regulating angular momentum, channelling accretion, and launching jets and outflows. While the magnetic properties of classical T Tauri stars (CTTSs) have been extensively characterised, those of their progenitors, Class I and flat-spectrum (FS) protostars, remain poorly constrained due to their embedded nature, which provides an observational challenge.
Aims. Our aim was to detect and characterise the large-scale magnetic fields in a sample of Class I and FS protostars. These young stars are expected to host strong magnetic fields generated by dynamo processes in their largely convective interiors.
Methods. We have used SPIRou, a high-resolution spectropolarimeter working in the near-infrared domain, to analyse the polarised light of Class I and FS protostars. We used the least-squares deconvolution (LSD) technique to perform the magnetic analysis and measure the longitudinal magnetic fields from circularly polarised Stokes $V$ profiles.
Results. We report new large-scale magnetic field detections in five FS protostars. Including the previous detection of the large-scale magnetic-field in the V347 Aur FS-protostar, 40% of our final sample of 15 protostars is confirmed to be magnetic. These magnetic stars show clear Zeeman signatures, with longitudinal field strengths ranging from $\sim$80 to $\sim$200 G in absolute value. The remaining stars exhibit no detectable Stokes $V$ signature, but the estimated upper limits on a hidden dipolar field range from 500 G to more than 5 kG. For stars in which no magnetic fields are detected, it is still conceivable that a magnetic field exists, but is intrinsically weak, highly complex and dominated by small-scale structures, or cancelled out in integrated spectropolarimetric signals due to opposing polarities.
Conclusions. We show that Class I and FS protostars can host large-scale magnetic fields with strengths that may be weaker than in more evolved CTTSs. This supports the idea that magnetic processes are already active during the main accretion phase and may influence star–disk interactions from the earliest stages.
@@ 1 Introduction
Star formation occurs in dense cores within molecular clouds. These clouds are magnetised, and the cores, called pre-stellar cores (Ward-Thompson et al., 1994), collapse under gravity, eventually giving birth to protostars. During this early evolution, magnetic fields are thought to play a significant role in several processes, such as accretion, ejection, and interaction with the circumstellar environment (Lada, 1987). Through their role in these processes, which significantly impact stellar rotation and evolution, magnetic fields shape the close environment of the protostar (Bouvier et al., 2007; Vaytet et al., 2018).
@@
Regulation of angular momentum is made through magnetospheric accretion and ejection processes during the pre-main sequence (PMS) phase (Bouvier et al., 2014; Lee et al., 2017). In classical T Tauri stars (CTTSs), the presence of large-scale magnetic fields is well established (Donati et al., 2011). These fields, typically of kilogauss strength, are responsible for truncating the inner disk and channelling the accreted material onto the stellar surface (Bouvier et al., 2007, 2014; Hartmann et al., 2016). The magnetic properties of CTTSs (between 0.5 and 1.3 $M_{\odot}$) are characterised by a dominant poloidal field, with an average strength between 1 and 2 kilogauss (Zaire et al., 2024).
@@
While magnetism and its impact on the star–disk interaction have been intensely explored during the PMS phase (Romanova et al., 2015), our current knowledge of the magnetic properties and of the star–disk interaction in earlier phases remains limited. During those phases, the protostar experiences intense accretion, often through bursts, and grows in mass until it approaches its final mass (Baraffe et al., 2009; Fischer et al., 2023). Such major internal structural changes may have a strong impact on the stellar magnetic field, which in turn may influence the connection between the star and its accretion disk (Kunitomo et al., 2017). The progenitors of the CTTSs (Class II) are the Class I protostars. These objects are still embedded within their envelopes of gas and dust (Lada, 1987). The protostars transiting from Class I to Class II are observed as flat-spectrum (FS) sources (Greene et al., 1994). Due to their dense and obscuring environments, the direct observation of Class I protostars and the detection of photospheric features is challenging (Connelley and Greene, 2010; Doppmann et al., 2005), making it difficult to study their surfaces, and therefore their magnetic fields.
@@
For several reasons we expect Class I protostars to be strongly magnetised, as in CTTSs. They have evolved enough to set off convection in their full interior (Palla, 1998; Baraffe et al., 2017), and are thus capable of hosting strong magnetic fields generated by dynamo processes. In other types of fully convective stars, such as the M dwarfs on the main sequence, strong dynamo magnetic fields are routinely observed (Donati et al., 2008; Morin et al., 2010; Donati et al., 2023). Signs of magnetospheric accretion, as in the CTTSs, have been detected in Class I sources (Doppmann et al., 2005; Fiorellino et al., 2023), suggesting the presence of strong large-scale magnetic fields at the surface of the embedded star. Powerful flares have been detected in X-ray observations of Class 0 and Class I protostars (Grosso et al., 1997, 2020; Montmerle et al., 2000), similar to those present in the Sun or more evolved stars, but much more energetic (Feigelson et al., 2007). These emissions are clues of intense magnetic activity. These studies therefore suggest that strong magnetic fields may already be in place during the Class I phase and may regulate mass inflow and angular momentum loss early on.
@@
A recent study by Flores et al. (2024) has investigated a sample of 42 Class I and FS protostars, and measured their total surface magnetic field from the Zeeman broadening of spectral lines. They found a median magnetic field strength of 1.9 kG, which is consistent with the value found in Class II sources and with earlier studies of Class I protostars (Johns-Krull et al., 2009), indicating that Class I and FS sources are indeed strongly magnetic. However, this study cannot inform us on the large-scale components of their magnetic fields, which interact with the disk. To understand whether magnetospheric accretion and ejection processes are in place at that age, and are potentially at the origin of the jets observed in almost all Class I sources, we need to characterise their large-scale magnetic field.
@@
Furthermore, dynamo processes are not well understood, especially in fully convective stars. The imprint of the dynamo processes in stellar evolution is still an open question (Moss, 2003; Donati et al., 2008). In particular, simulations have shown that the initial conditions at the launch of the dynamo may impact the future properties of dynamo magnetic fields, and may explain the bistability of magnetic fields observed in fully convective M dwarfs (Gastine et al., 2012, 2013; Raynaud et al., 2015). While the magnetic energy produced by dynamo processes is an important characteristic to constrain the models, another factor is the ratio of poloidal to toroidal fields, which is only accessible by measuring the magnetic geometry.
@@
The topology and strength of the large-scale magnetic field can be obtained with the Zeeman Doppler imaging (ZDI) technique (Semel, 1989). This technique uses time series of circularly polarised and intensity spectra well sampled over the rotation period of the star using a high spectral resolution spectropolarimeter. By assuming that the magnetic field components can be modelled as spherical harmonics, and by using an inversion technique, ZDI provides us with the maps of the radial, azimuthal, and meridional components of the field (Donati et al., 2006; Kochukhov, 2016). This technique has been largely employed to map the magnetic field of various types of stars, including CTTSs (e.g. Donati and Gregory, 2009; Zaire et al., 2024).To date, ZDI has not been largely performed in Class I and FS protostars.
@@
Recent advances in instrumentation have opened new possibilities for investigating these elusive protostars. Among these instruments is the SPectropolarimètre InfraRouge (SPIRou), a spectropolarimeter installed on the 3.6-meter Canada-France-Hawaii Telescope (CFHT). SPIRou (Donati et al., 2020) operates in the near-infrared (nIR) domain, which allows us to pierce through their dusty envelope to probe the photosphere of Class I protostars. SPIRou’s high-resolution, high-sensitivity measurements in the $YJHK$ bands are particularly well suited to study Class I protostars, which are largely inaccessible to optical spectropolarimeters.
@@
Before performing a monitoring for each source in our sample of Class I and FS protostars, to map their magnetic fields, we performed a snapshot survey with SPIRou. Our aim was to first detect their Zeeman signature in circularly polarised (Stokes $V$) spectra. In future work, we will monitor with SPIRou only those with a Zeeman detection to perform ZDI.
In this study, we present the analysis of our snapshot sample. It is structured as follows. We describe the observations, data reduction, and spectral and polarimetric analyses in Sect. 2. Our results are presented in Sect. 3 and discussed in Sect. 4. We summarise our conclusions in Sect. 5.
@@ 2.1 Sample selection
Our goal is to characterise the magnetic fields of a sample of Class I and FS objects. To this aim, we need to detect the circular polarisation signal induced by the Zeeman effect in the photospheric lines. To build our sample, we first listed the objects in which photospheric lines had been detected in previous surveys using high-resolution optical or nIR spectrographs (White and Hillenbrand, 2004; Doppmann et al., 2005; Evans et al., 2009; Connelley and Greene, 2010; Rebull et al., 2010; Viana Almeida et al., 2012; Dunham et al., 2015). Then we first kept the sources north to a declination of -40, to ensure that they were observable from the CFHT. In a second step, we selected sources brighter than 10 magnitudes in the H band so that the total exposure time needed to detect a typical dipolar magnetic field of $\sim$0.5 kilogauss is within 1.5 h, assuming a typical veiling in the H band of about 2 (Doppmann et al., 2005). In the star-forming regions that we explored, Class I and FS sources display H band magnitudes typically ranging from 8 to 12, depending on how deeply embedded they are or how inclined their bipolar cavity is with respect to our line of sight. As a consequence, the magnitude criterion may introduce a bias towards the less embedded objects, i.e. more evolved Class I objects, or towards low inclination. Two of our targets have magnitudes H greater than 10. We had an opportunity to obtain additional observations, so we included them in our sample.
@@ 2.2 SPIRou observations and data reduction
The primary goal of SPIRou is to detect and characterise exoplanets around M dwarfs, but also to study magnetic fields in low-mass mature and PMS (Donati et al., 2020). SPIRou covers a spectral range from 950 nm to 2 500 nm, with a resolution of 70 000 and a velocity bin per pixel of 2.3 km s-1(Donati et al., 2020). It consists of an achromatic polarimeter, which splits light into two beams associated with orthogonal states of the selected polarisation, transmitted via optical fibres to an ultra-stable thermally controlled cryogenic spectrograph. The spectrograph disperses light using an echelle grating and a double-pass cross-dispersing prism train, recording the spectra on an H4RG detector.
The polarimeter consists of Fresnel rhombs and a Wollaston prism, enabling simultaneous measurements of both orthogonal states of the selected polarisation (circular polarisation in the present case). The operational process includes stabilising the stellar image on the entrance aperture, injecting it into optical fibres, and using calibration sources (a Fabry-Pérot interferometer, in our case) for precise wavelength and radial velocity calibration.
@@
We used V0.7.292 version of the dedicated APERO [^2] https://apero.exoplanets.ca/ pipeline (Cook et al., 2022) to reduce the observations. Compared to version V0.7.255 described in Cook et al. (2022), in the APERO versions there have been changes to the background subtraction. The thermal background spectrum is obtained by averaging the spectra acquired from telescope dark frames (i.e. with no sky scene injected). The thermal spectrum is then scaled to match the thermal emission of saturated absorption lines in the K band. An issue was noted, as persistence structures in the J band led to spurious subtraction, being mistaken for low-level thermal emission. Since one does not expect a measurable level of thermal emission below 1.9 µm (considering the rapid decline of a near-room-temperature blackbody bluewards of this wavelength), the thermal emission spectrum was set to zero below 1.9 µm in the newest versions of APERO. It should be noted that, in practise, the thermal background is only subtracted within the K band; thus, this zeroing of the thermal emission at the blue end does not affect the overall thermal background correction. Finally, all individual spectra were corrected for telluric absorption (including OH lines), as described in Cook et al. (2022).
@@ 2.3 Continuum normalisation of individual spectra
To analyse the magnetic signals within the photospheric lines, a normalisation of the spectra to their continuum is needed. Automatic continuum normalisation can be performed during the extraction of the polarised spectra. However, the spectra of Class I sources have several variable emission lines that need to be taken into account on a case-by-case basis. While normalisation tools already exist, they lack specific options we aimed to incorporate (necessary for emission line spectra such as those of Class I sources). We developed a normalisation tool for SPIRou data files called Fantasio. [^3] https://github.com/Lisa2626/Fantasio The code aims to reproduce the principles of the IRAF normalisation tool called continuum. The selection of the continuum points is performed by removing non-continuum points using a sigma-clipping procedure over several iterations. The normalisation is performed order by order when a 2D SPIRou spectrum is given as input. The developed interface allows users to modify normalisation parameters interactively for each order, including selecting the polynomial order of the spline and the number of spline interpolations. The sigma-clipping parameters, such as the upper and lower thresholds, and the number of iterations, can also be adjusted to optimise the fit. We allow different threshold values for points above and below the continuum to provide flexibility when both photospheric and emission lines are present within one order. Another interactive feature enables the removal of photospheric and emission lines to exclude them from the fitting process by selecting the regions to reject. This is particularly useful when dealing with spectra containing numerous strong emission lines, as in Class I sources. Normalisation can be performed manually by selecting parameters for each order, or it can be run automatically for multiple observations using pre-defined parameters.
Figure 1: Upper panel: Order 32 of the spectrum of VV CrA SW without normalisation (solid black line). In blue are the data points selected to fit the continuum using a sigma-clipping of 3 and 2 iterations. The spline fit of order 3 is plotted in red with a number of knots of 2. The two regions with emission lines were manually removed using the interactive tool. Lower panel: Same spectrum after normalisation. The dashed red line shows the continuum normalised to 1.
We performed a manual normalisation on both A- and B-spectra, corresponding to the two orthogonal polarisations, for one observation of each star. We then applied the automatic normalisation to all other observations of the same star, using the same fitting parameters per order as those chosen for the fit of the first observation. In general, the order of the spline fit ranges from 1 to 3, as the number of knots of the spline. The sigma clipping can vary from 1 to 3 sigma, and 3 iterations are used. An example of the normalisation process is shown in Fig. 1.